# Style Transfer: From Stitching to Neural Networks

### v2

@@ I Introduction
The introduction of deep learning to style transfer has notably enhanced the quality and flexibility of the transformations. Utilizing convolutional neural networks (CNNs) [22, 23, 24, 25, 26], deep-learning-based approaches analyze and abstract different levels of features in low-level features for basic textures and edges, and high-level features for more complex structures and semantics. In this framework, a pre-trained network like VGG [27, 28] is often used to extract these features. This method not only provides stunning visual results but also offers the flexibility to blend multiple styles and control the extent of stylization effectively.
@@ II-A Traditional Style Transfer Method
Traditional style transfer methods in computer vision focus on reimagining an image by blending its content with the stylistic elements of another image. These techniques primarily involve non-neural approaches such as texture synthesis, patch-based methods, and optimization-based techniques [29, 10, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Some methods utilize 3D information [40, 41, 42, 43, 44, 45], while others rely solely on plain image data [10, 46, 47, 48, 49]. In this paper, we select one of the most well-known texture methods by Efros et al. [10] for the comparative study.
@@ II-B Deep-learning-based Style Transfer
Deep-learning based methods in computer vision have revolutionized style transfer by leveraging neural networks to blend the content of one image with the stylistic elements of another. These techniques predominantly involve CNNs [23, 24, 50, 51, 52, 53, 54, 55, 56] and many other machine learning techniques [57, 58, 24, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], which learn to synthesize new images through deep representations. Unlike traditional methods, deep-learning approaches can capture complex patterns and textures, offering more flexibility and realism in the generated images. In this paper, we explore one of the most representative deep-learning style transfer methods introduced by Ding et al., which utilizes a novel segmentation framework to achieve impressive results in artistic image synthesis [73].
@@ IV-B Future Work
While this study provides a thorough analysis comparing traditional and deep-learning-based style transfer methods, several limitations could serve as avenues for future research. First of all, the cost-effectiveness of both traditional and deep-learning-based methods deserves further examination. Understanding the computational and resource requirements of each approach will be crucial for evaluating their practicality in real-world applications. Secondly, additional studies incorporating a wider range of diverse strategies should be undertaken to enable a more comprehensive discussion of their strengths and weaknesses. Exploring hybrid techniques that combine elements from both paradigms could lead to innovative solutions that better capture artistic styles [75, 76].

### v3

@@ I Introduction
The introduction of deep learning to style transfer has notably enhanced the quality and flexibility of the transformations. Utilizing convolutional neural networks (CNNs) [22, 23, 24, 25], deep-learning-based approaches analyze and abstract different levels of features in low-level features for basic textures and edges, and high-level features for more complex structures and semantics. In this framework, a pre-trained network like VGG [26, 27] is often used to extract these features. This method not only provides stunning visual results but also offers the flexibility to blend multiple styles and control the extent of stylization effectively.
@@ II-A Traditional Style Transfer Method
Traditional style transfer methods in computer vision focus on reimagining an image by blending its content with the stylistic elements of another image. These techniques primarily involve non-neural approaches such as texture synthesis, patch-based methods, and optimization-based techniques [28, 10]. Some methods utilize 3D information [29, 30], while others rely solely on plain image data [10, 31, 32, 33]. In this paper, we select one of the most well-known texture methods by Efros et al. [10] for the comparative study.
@@ II-B Deep-learning-based Style Transfer
Deep-learning based methods in computer vision have revolutionized style transfer by leveraging neural networks to blend the content of one image with the stylistic elements of another. These techniques predominantly involve CNNs [23, 24, 34] and many other machine learning techniques [35, 24, 36], which learn to synthesize new images through deep representations. Unlike traditional methods, deep-learning approaches can capture complex patterns and textures, offering more flexibility and realism in the generated images. In this paper, we explore one of the most representative deep-learning style transfer methods introduced by Ding et al., which utilizes a novel segmentation framework to achieve impressive results in artistic image synthesis [37].
@@ IV-B Future Work
While this study provides a thorough analysis comparing traditional and deep-learning-based style transfer methods, several limitations could serve as avenues for future research. First of all, the cost-effectiveness of both traditional and deep-learning-based methods deserves further examination. Understanding the computational and resource requirements of each approach will be crucial for evaluating their practicality in real-world applications. Secondly, additional studies incorporating a wider range of diverse strategies should be undertaken to enable a more comprehensive discussion of their strengths and weaknesses. Exploring hybrid techniques that combine elements from both paradigms could lead to innovative solutions that better capture artistic styles [39].